OpenStudy (anonymous):

What is equation in point slope form for a line with a point at (4,-5) and a slope of 2? THANK YOU!!

OpenStudy (anonymous):

\[y+5=2(x-4)\] Then you solve it out to put it in y=mx+b form \[y=2x-13\]

OpenStudy (ujjwal):

Use relation y-y1=m(x-x1) where (x1, y1) is given point and m is slope

OpenStudy (anonymous):

Thank you guys:) I was having a hard time this.